Exercise 14-04

On October 31, the stockholders’ equity section of Cullumber Company consists of common stock $260,000 and retained earnings of $882,000. Cullumber is considering the following two courses of action: (1) declaring a 4% stock dividend on the 26,000, $10 par value shares outstanding, or (2) affecting a 2-for-1 stock split that will reduce par value to $5 per share. The current market price is $16 per share.

Prepare a tabular summary of the effects of the alternative actions on the components of stockholders’ equity, outstanding shares, and par value per share.




Before
Action


After
Stock
Dividend


After
Stock
Split

Stockholders’ equity





   Paid-in capital





      Common stock
$


$


$

      In excess of par



         Total paid-in capital



   Retained earnings



         Total stockholders’ equity
$


$


$

Outstanding shares



Par value per share
$


$


$
Answer #1


Prepare a tabular summary of the effects of the alternative actions on the components of stockholders’ equity, outstanding shares, and par value per share.


Before
Action

After
Stock
Dividend

After
Stock
Split

Stockholders’ equity
   Paid-in capital
      Common stock $260000

$27040*10 = 270400

$

260000

      In excess of par

26000*4%*6 = 6240

0

         Total paid-in capital

260000

276640

260000

   Retained earnings

882000

882000-1040*16 = 865360

882000

         Total stockholders’ equity $1142000

$1142000

$1142000

Outstanding shares

26000

26000*1.04 = 27040

26000*2 = 52000

Par value per share $10

$10

$10/2 = 5